Generation Next has Tk 2000 million of total asset of which Tk 990 million is fixed asset and the rest is current asset. 20% of currents assets are needed for temporary seasonal need and the rest is required year-round. If the company follows hedging approach in financing its working capital how much of long-term and short-term borrowing will the company have


1
Expert's answer
2021-12-09T10:19:21-0500

Maturity matching or hedging approach is a strategy of working capital financing wherein we finance short term requirements with short-term debts and long-term requirements with long-term debts.

Long Term Funds will Finance = Fixed Assets + Permanent Working Capital

Short Term Funds will Finance = Temporary Working Capital


current assets=2000-990=1010

Permanent Working Capital=1010-202=808

"temporary seasonal need=0.2\times1010=202"


Long Term Funds will Finance = Fixed Assets + Permanent Working Capital=990+808=1798

Short Term Funds will Finance = Temporary Working Capital=202